October 19, 2016 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, Septemeber 21, 2016
11:00 am, EST
JLab: CEBAF Center, F326

Communication Information

Remote Connection

You can connect using BlueJeans using the meeting number 630 804 895 .       (Click "Expand" to the right for more details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 630804895
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or bjn.vc
    • Enter meeting ID above
    • Use *4 to unmute


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2016/ .

Calibration Tasks

  1. Spring 2016 Recon Checklist
  2. Mcsmear updates - mcsmear 20160323
  3. Data Validation
  4. GlueX Detector Performance Document

Upcoming Calibrations

  • Updated error matrices
    • BCAL (Mark D.)
    • FCAL (Mark D.)
    • Tracking (Simon)
  • FCAL timing updates (Adesh)
  • Add timing information to FCAL clustering/inner ring calibrations? (Adesh/Matt S.)
  • BCAL TDC calibrations (Andrei)
  • FDC per-channel calibrations (Alex A./Mike S.)
  • TAGH alignment (Nathan)
  • Tagger energy map (Richard J.?)
  • TOF alignment (Beni/Simon)

New fADC250 Data Format Offline tasks

  • ✓ EVIO Parsing (David L.) [waiting to test with data]
  • ✓ Calibrated Hit Creation (Sean) [waiting to test with data]
    • ✓ Low-level monitoring plugin (Sean)
  • Update monitoring plugins
    • ✓ plugins/monitoring/BCAL_online/ (Mark D.)
    • ✓ plugins/monitoring/BCAL_LEDonline/ (Ahmed)
    • ✓ plugins/monitoring/DAQ_online/ (David L.?)
    • ✓ plugins/monitoring/FCAL_online/ (Adesh)
    • ✓ plugins/monitoring/highlevel_online/ (Paul M.)
    • ✓ plugins/monitoring/L1_online/ (Alex S.)
    • ✓ plugins/monitoring/pedestal_online/ (?)
    • ✓ plugins/monitoring/PS_online/ (Nathan)
    • ✓ plugins/monitoring/PSC_online/ (Nathan)
    • ✓ plugins/monitoring/ST_online_lowlevel/ (Mahmoud)
    • ✓ plugins/monitoring/TAGH_online/ (Nathan)
    • ✓ plugins/monitoring/TPOL_online/ (Nathan)
  • Update calibration
    • plugins/Calibration/TOF_calib/ (Beni)
    • ✓ plugins/Calibration/st_tw_corr_auto/ (Mahmoud)
  • Update other plugins
    • ✓ plugins/Analysis/pedestals/ (?)
    • plugins/Analysis/DAQTree/ (?)
    • plugins/Analysis/DAQTreeBCAL/ (?)
    • ✓ plugins/Utilities/evio_writer/ (Sean/David L.) [waiting to test]


  1. Announcements
  2. Reconstruction Launch Readiness
  3. Run Planning
  4. Calibration Updates
  5. Simulations
    • Data/Simulation comparisons
  6. AOB


Attending: Sean, Luke (NU); Simon, Adesh, Valery L., Brad, Thomas, Alex A., Eugene, Mark I., Sergey (JLab); Curtis, Mike S., Will M., Naomi (CMU); Matt S. (IU); Alex B., Paul, Beni, Mark D. (BlueJeans land)

  1. Announcements
    • Mark I. announced that there is a test ifarm node running CentOS 7, and that the GlueX software stack has been successfully compiled. Sean ran monitoring_hists over one file, and noticed some tracking problems (reconstruction issue tracking see GitHub issue).
  2. Reconstruction Launch Readiness
    • Paul reported that the monitoring launch started last Friday successfully finished, and there will be one more monitoring launch before the reconstruction launch to make sure the reconstruction looks good.
    • New items: Adesh added new FCAL timing calibrations on Monday; Mark Dalton is making some small changes to the BCAL/FCAL covariance matrices.
    • Mike S. has the code to perform the updated FDC timing calibrations, he's currently looking at the results for several runs to finalize the procedure for the whole run period.
    • A couple other issues were recently brough up. Mahmoud noticed that there is a hole in the SC/track matching, and during the meeting we looked at some real-time results (see GitHub issue). Alex B. pointed out this morning that the timing for some TAGM channels seems to be wrong and he is looking into it.
  3. Run Planning
    • New fADC250 data format As of yesterday, we have cosmic data with the BOR/config information correctly propagated. Sean has been checking the new fADC250 firmware by comparing the reported pulse information and to the emulated information, and has noticed a few problems including a potentially reoccuring problem with integration windows. Follow the sim-recon sdobbs_fadc250emu_updates branch for more developments on this front.
    • Naomi has started looking at the fADC125 data without much luck. Beni has been looking at CDC waveforms, and everything looks good except for ~8 channels.
    • Mark Dalton has been looking at the BCAL data and working through some issues.
    • Adesh updated the FCAL monitoring plugins & his calibration scripts for ROOT 6. So far the new data looks good, except for one bad fADC board.
    • We reviewed ROOT 6 progress, and are most of the way there.
    • Sean will organize a test calibration run over some of the cosmic data before the next meeting and work on debugging the workflow.
  4. Calibration Updates
    • FCAL - Adesh has been tuning up his monitoring and calibration scripts to work with the new data. He is looking at updating the MC time resolution parameters.
    • BCAL - Besides the small updates to the covariance matrices, Mark Dalton has been looking at improving the BCAL timing. When he properly calibrates the TDCs, he is not seeing much improvement in shower time resolution, so he is revisiting how we calculate the shower times. Will M. found that some of the time smearing in mcsmear was being done incorrectly, so he fixed this and is revisiting his data/MC comparisons.
    • FDC - Lubomir doesn't expect much run-to-run variation in the wire time calibrations. Mike S. is finding a change in calibration from what is currently in the CCDB, which increases the number of reconstructed tracks by a significant amount.
    • TOF - Beni reported that Ed has shown that we can load new non-linear compensation lookup tables and get results that are consistent with doing the corrections offline, using the firmware updates from Bryan Moffit. Beni is collecting more statistics and plans to implement this new calibration on Monday. His ROOT6 issues are mostly resolved, although he is getting some residual differences in calibrations.
    • TAGM - When Alex B. was updating the timewalk corrections with improved values, he noticed that the calibrations for some channels in some runs had changed. He is looking into this.
    • TAGH - Nathan has updated his calibration scripts and is ready for the upcoming run.
  5. Simulations
    • See above for discussions. Simon asked Mark D. about the updated calorimeter covariance matrices, and he confirmed that they were determined using simulations, so they are applicable to data and simulation, and will be updated when Will updates the BCAL simulations.